# Milky Way-est: Cosmological Zoom-in Simulations with Large Magellanic Cloud and Gaia–Sausage–Enceladus Analogs

### v1

@@ 5.1.2 Observational Implications
Our predictions are also relevant for interpreting satellite galaxy populations beyond the MW. In particular, recent surveys have characterized satellite populations of MW-mass centrals throughout the Local Volume (e.g., ELVES; Carlsten et al. 2022) and at distances of $25$ to $40~{}{\rm Mpc}$ (e.g., SAGA; Geha et al. 2017; Mao et al. 2021; Y.-Y. Mao et al., in preparation). Interestingly, SAGA satellite abundances correlate strongly with brightest satellite luminosity, which is reminiscent of the strong impact that Milky Way-est LMC analogs have on total subhalo abundance. Thus, we expect that Milky Way-est simulations will inform future comparisons between MW, Local Volume, and SAGA satellite populations.
Beyond satellite galaxy modeling, we anticipate that Milky Way-est simulations will help refine predictions for the MW’s stellar stream population, and particularly for the subhalos that perturb stellar streams. Current stream observations are most sensitive to subhalos in the inner regions of the MW (e.g., Bonaca et al. 2019; Banik et al. 2021). Thus, the abundant and centrally-concentrated nature of Milky Way-est subhalo populations relative to typical MW-mass hosts has important implications for stream modeling. Spatial anisotropy in the distribution of stream perturbers over the full sky may also become detectable as stream observations improve (Drlica-Wagner et al., 2019). Resimulations of Milky Way-est systems at higher resolution will resolve even lower-mass subhalos associated with the LMC, which will be relevant for interpreting upcoming stellar stream measurements (e.g., Arora et al. 2023).
@@ 5.2 Simultaneously Modeling the Impact of GSE and LMC
First, Milky Way-est hosts’ specific dark matter accretion rates are enhanced during the GSE merger, consistent with previous simulation results (e.g., Evans et al. 2020). However, during this epoch, Milky Way-est hosts’ masses are lower than those of average hosts with the same final mass due to the LMC’s large contribution to Milky Way-est hosts’ total masses at late times. For example, the mean mass of Milky Way-est hosts is $1.0\times 10^{11}~{}{M_{\odot}}$ at $a=0.25$, versus $1.6\times 10^{11}~{}{M_{\odot}}$ in Symphony MW. Thus, it is timely to revisit the impact of GSE on disk and stellar halo formation in low-mass MW progenitors.
Second, various studies have shown that only a small fraction ($\approx 10\%$) of Milky Way-mass halos in cosmological simulations host LMC and SMC-mass subhalos (e.g., Boylan-Kolchin et al. 2010; Busha et al. 2011; Liu et al. 2011). We find that an even smaller fraction ($<1\%$) of hosts have both recently-accreted, nearby LMCs and GSE analogs, where the LMC’s present-day distance is the most constraining cut (see Table 1). Qualitatively, this result is not surprising; adding constraints makes the resulting sample rarer. Practically, the number density of hosts that satisfy our criteria is $33$ systems in a $(125~{}{\rm Mpc}~{}h^{-1})^{3}$ volume. Thus, simulating a large number of halos with both GSE and LMC analogs at sufficient resolution to study their low-mass subhalo populations is challenging. Future work directly generating MW-like hosts with LMC and GSE analogs, and perhaps additional systems like SMC analogs, will be useful. Constrained simulations (e.g., Rey et al. 2019) and semi-analytic merger tree realizations (e.g., Nadler et al. 2023a) are promising in this regard.
@@ 5.3 Caveats
We now discuss several caveats. First, our simulations do not include baryons, which are known to affect subhalo populations, e.g. via enhanced disruption of subhalos that orbit near the central galaxy (e.g., D’Onghia & Lake 2008; Garrison-Kimmel et al. 2017; Kelley et al. 2019). Unlike existing zoom-in simulations where this disk potential technique has been applied, Milky Way-est systems are explicitly selected to accrete both GSE and LMC analogs. Incorporating disk potentials in resimulations of Milky Way-est systems will be a step toward modeling the impact of baryons on their subhalo populations. Because the disk preferentially disrupts subhalos that accrete early (Garrison-Kimmel et al., 2017; Nadler et al., 2018), including central galaxy potentials will likely enhance the contribution of recently-accreted LMC-associated subhalos to host subhalo populations, particularly in the inner regions, while further suppressing the contribution of GSE-associated subhalos.
Second, we have defined “MW-like” halos in our Milky Way-est suite based on halo mass, concentration, and the existence of GSE and LMC analogs. Even with all of these criteria, other features of Milky Way-est systems may not be representative of the MW; for example, the large-scale environment—which is not explicitly captured by the LMC and GSE constraints we impose—may influence host halo formation histories (e.g., Garrison-Kimmel et al. 2019; Santistevan et al. 2020). Furthermore, it is unclear whether Milky Way-est hosts form galaxies consistent with the MW galaxy’s observed properties. We plan to apply empirical (e.g., UniverseMachine; Behroozi et al. 2019; Wang et al. 2021) and semi-analytic (e.g., GRUMPY; Kravtsov & Manwadkar 2022) models to gain insight into galaxy formation in Milky Way-est systems. Milky Way-est systems are excellent candidates for zoom-in resimulation with full hydrodynamic and galaxy formation physics.
Finally, our analysis may underestimate the contribution of GSE-associated subhalos. In particular, although Rockstar and consistent-trees perform well compared to other subhalo finders and merger tree algorithms (e.g., Onions et al. 2012; Srisawat et al. 2013), these tools can lose track of early-accreted subhalos in a nonphysical manner (Mansfield et al., 2023). It will be important to study GSE analogs and their contribution to present-day Milky Way-est subhalo populations using improved halo catalogs and merger trees based on, e.g., particle tracking.
@@ 6 Conclusions
We have presented Milky Way-est, a suite of 20 cosmological cold-dark-matter-only zoom-in simulations of Milky Way analogs selected to undergo LMC and GSE-like merger events. By comparing Milky Way-est with hosts that are not specifically selected to undergo these mergers, both from their parent cosmological volume and from the Symphony MW zoom-in simulations (Nadler et al., 2023b), we find that:
@@
It is interesting to consider how many of (and how accurately) the MW’s major accretion events must be modeled to meet the precision of upcoming MW dwarf galaxy and stellar stream observations. On even smaller, sub-kiloparsec scales, stellar velocity distributions near the solar circle and dark matter direct detection experiments are sensitive to substructure accreted during specific merger and accretion events like the GSE and LMC (e.g., Necib et al. 2019; Besla et al. 2019; Smith-Orlik et al. 2023). This motivates studies of the local dark matter distribution using Milky Way-est.
@@ Acknowledgements
We plan to release Milky Way-est halo catalogs and merger trees upon publication. We are grateful to Elise Darragh-Ford, Nitya Kallivayalil, Philip Mansfield, Yunchong Wang, and Andrew Wetzel for comments on the manuscript and to Susmita Adhikari and Arka Banerjee for helpful conversations related to this work.
This research was supported in part by the Kavli Institute for Particle Astrophysics and Cosmology (KIPAC) at Stanford University and SLAC National Accelerator Laboratory and by the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory. D.B. received additional support from the Stanford University Summer Undergraduate Research Program in Physics. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
This research made use of computational resources at SLAC National Accelerator Laboratory, a U.S. Department of Energy Office of Science laboratory and the Sherlock cluster at the Stanford Research Computing Center (SRCC); the authors are thankful for the support of the SLAC and SRCC computational teams. This work used the Extreme Science and Engineering Discovery Environment (XSEDE) Stampede2 and Ranch systems at the Texas Advanced Computing Center (TACC) through allocation TG-PHY200009. The authors acknowledge TACC at The University of Texas at Austin for providing HPC resources that have contributed to the research results reported within this paper.
@@ Appendix B Convergence Test
Figure 10: SHMF for Milky Way-est Halo004 at our fiducial resolution (purple) compared to a higher-resolution resimulation (gold). For the fiducial-resolution case, we use our standard subhalo mass cut of $M_{\mathrm{sub}}>300\times m_{\mathrm{part}}=1.2\times 10^{8}~{}{M_{\odot}}$; for the higher-resolution case, we use $M_{\mathrm{sub}}>300\times m_{\mathrm{part,high-res}}=1.5\times 10^{7}~{}{M_{\odot}}$. The bottom panel shows the ratio of fiducial to higher-resolution measurements, along with ±$\pm 10\%$ deviations (thin dotted lines) about a ratio of one (dotted line).

### v2

@@ 5.1.2 Observational Implications
[…]
Beyond satellite galaxy modeling, we anticipate that Milky Way-est simulations will help refine predictions for the MW’s stellar stream population, particularly for the subhalos that perturb stellar streams. Current stream observations are most sensitive to subhalos in the inner regions of the MW (e.g., Bonaca et al. 2019; Banik et al. 2021). Thus, the abundant and centrally concentrated nature of Milky Way-est subhalo populations relative to typical MW-mass hosts has important implications for stream modeling. Spatial anisotropy in the distribution of stream perturbers over the full sky may also become detectable as stream observations improve (Drlica-Wagner et al., 2019). Resimulations of Milky Way-est systems at higher resolution will resolve even lower-mass subhalos associated with the LMC, which will be relevant for interpreting upcoming stellar stream measurements (e.g., Arora et al. 2023).
@@ 5.2 Simultaneously Modeling the Impact of GSE and LMC
First, Milky Way-est hosts’ specific dark matter accretion rates are enhanced during the GSE merger, consistent with previous simulation results (e.g., Evans et al. 2020). However, during this epoch, Milky Way-est hosts’ masses are lower than average hosts with the same final mass due to the LMC’s large contribution to Milky Way-est hosts’ total masses at late times. For example, the mean mass of Milky Way-est hosts is $1.0\times 10^{11}~{}{M_{\odot}}$ at $a=0.25$ ($z=3$), versus $1.6\times 10^{11}~{}{M_{\odot}}$ in Symphony MW. Thus, it is timely to revisit the impact of GSE on disk and stellar halo formation in low-mass MW progenitors.
Second, various studies have shown that only a small fraction ($\approx 10\%$) of Milky Way-mass halos in cosmological simulations host LMC- and SMC-mass subhalos (e.g., Boylan-Kolchin et al. 2010; Busha et al. 2011; Liu et al. 2011). We find that an even smaller fraction ($<1\%$) of hosts have both recently accreted, nearby LMCs and GSE analogs, where the LMC’s present-day distance is the most constraining cut (see Table 1). Qualitatively, this result is not surprising: adding constraints makes the resulting sample rarer. Practically, the number density of hosts that satisfy our criteria is $33$ systems in a $(125~{}{\rm Mpc}~{}h^{-1})^{3}$ volume. Thus, simulating a large number of halos with both GSE and LMC analogs at sufficient resolution to study their low-mass subhalo populations is challenging. Future work directly generating MW-like hosts with LMC and GSE analogs and perhaps additional systems like SMC analogs will be useful. Constrained simulations (e.g., Rey et al. 2019) and semi-analytic merger-tree realizations (e.g., Nadler et al. 2023a) are promising in this regard.
@@ 5.3 Caveats
We now discuss several caveats. First, our simulations do not include baryons, which are known to affect subhalo populations, e.g. via enhanced disruption of subhalos that orbit near the central galaxy (e.g., D’Onghia & Lake 2008; Garrison-Kimmel et al. 2017; Kelley et al. 2019). Unlike existing zoom-in simulations where this disk potential technique has been applied, Milky Way-est systems are explicitly selected to accrete both GSE and LMC analogs. Incorporating disk potentials in resimulations of Milky Way-est systems will be a step toward modeling the impact of baryons on their subhalo populations. Because the disk preferentially disrupts subhalos that accrete early (Garrison-Kimmel et al., 2017; Nadler et al., 2018), including central galaxy potentials will likely enhance the contribution of recently accreted LMC-associated subhalos to host subhalo populations, particularly in the inner regions, while further suppressing the contribution of GSE-associated subhalos.
Second, we have defined “MW-like” halos in our Milky Way-est suite based on halo mass, concentration, and the existence of GSE and LMC analogs. Even with all of these criteria, other features of Milky Way-est systems’ accretion histories may not be representative of the MW. For example, as noted in Section 2.3, we have not required Milky Way-est hosts to accrete a Sgr analog or remain quiescent between GSE disruption and the accretion of a potential Sgr analog; for Milky Way-est hosts with potential Sgr analogs, it will be interesting to explore how many Sgr-associated subhalos survive as the system tidally disrupts. Meanwhile, large-scale environment—which is not directly captured by our LMC and GSE constraints—may further influence host halo formation histories (e.g., Garrison-Kimmel et al. 2019; Santistevan et al. 2020). Beyond dark matter assembly histories, it is unclear whether Milky Way-est hosts form galaxies consistent with the MW galaxy’s observed properties. We plan to apply empirical (e.g., UniverseMachine; Behroozi et al. 2019; Wang et al. 2021, 2024) and semi-analytic (e.g., GRUMPY; Kravtsov & Manwadkar 2022) models to gain insight into galaxy formation in Milky Way-est systems. Milky Way-est systems are excellent candidates for zoom-in resimulation with hydrodynamic and galaxy formation physics.
Finally, our analysis may underestimate the contribution of GSE-associated subhalos. In particular, although Rockstar and consistent-trees perform well compared to other subhalo finders and merger-tree algorithms (e.g., Onions et al. 2012; Srisawat et al. 2013), these tools can lose track of early-accreted subhalos in a nonphysical manner (Mansfield et al., 2023). It will be important to study GSE analogs and their contribution to present-day Milky Way-est subhalo populations using improved halo catalogs and merger trees based on, for example, particle tracking.
@@ 6 Conclusions
We have presented Milky Way-est, a suite of 20 cosmological cold-dark-matter-only zoom-in simulations of Milky Way analogs selected to undergo LMC- and GSE-like merger events. By comparing Milky Way-est with hosts that are not specifically selected to undergo these mergers, both from their parent cosmological volume and from the Symphony MW zoom-in simulations (Nadler et al., 2023b), we find the following:
@@
It is interesting to consider how many of (and how accurately) the MW’s major accretion events must be modeled to meet the precision of upcoming MW dwarf galaxy and stellar stream observations. On even smaller, subkiloparsec scales, stellar velocity distributions near the solar circle and dark matter direct-detection experiments are sensitive to substructure accreted during specific merger and accretion events like the GSE and LMC (e.g., Necib et al. 2019; Besla et al. 2019; Smith-Orlik et al. 2023). This motivates studies of the local dark matter distribution using Milky Way-est.
@@ Acknowledgements
Milky Way-est data products are publicly available at http://web.stanford.edu/group/gfc/gfcsims/.
We thank the anonymous referee for their valuable feedback. We thank Matthew Becker for his work on the Chinchilla simulation suite used in this study. We are grateful to Elise Darragh-Ford, Nitya Kallivayalil, Philip Mansfield, Ekta Patel, Yunchong Wang, and Andrew Wetzel for comments on the manuscript and to Susmita Adhikari, Christian Aganze, and Arka Banerjee for helpful conversations related to this work.
This research was supported in part by the Kavli Institute for Particle Astrophysics and Cosmology (KIPAC) at Stanford University and SLAC National Accelerator Laboratory and by the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory. D.B. received additional support from the Stanford University Summer Undergraduate Research Program in Physics. This work was performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant No. PHY-2210452.
This research made use of computational resources at SLAC National Accelerator Laboratory, a U.S. Department of Energy Office of Science laboratory, and the Sherlock cluster at the Stanford Research Computing Center (SRCC); the authors are thankful for the support of the SLAC and SRCC computational teams. This work used the Extreme Science and Engineering Discovery Environment (XSEDE) Stampede2 and Ranch systems at the Texas Advanced Computing Center (TACC) through allocation TG-PHY200009. The authors acknowledge TACC at The University of Texas at Austin for providing HPC resources that have contributed to the research results reported within this paper.
This work used data from the Symphony suite of simulations (http://web.stanford.edu/group/gfc/symphony/), which was supported by the Kavli Institute for Particle Astrophysics and Cosmology at Stanford University and SLAC National Accelerator Laboratory, and by the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory.
@@ Appendix B Convergence Test
Figure 10: SHMF for Milky Way-est Halo004 at our fiducial resolution (purple) compared to a higher-resolution resimulation (gold). For the fiducial-resolution case, we use our standard subhalo mass cut of $M_{\mathrm{sub}}>300\times m_{\mathrm{part}}=1.2\times 10^{8}~{}{M_{\odot}}$; for the higher-resolution case, we use $M_{\mathrm{sub}}>300\times m_{\mathrm{part,high-res}}=1.5\times 10^{7}~{}{M_{\odot}}$. The bottom panel shows the ratio of fiducial to higher-resolution measurements, along with ±$\pm 10\%$ deviations (thin dotted lines) about a ratio of 1 (dotted line).